From acd@roe.ac.uk Thu Sep 15 12:22:36 2016 Date: Thu, 15 Sep 2016 11:22:26 +0000 From: Clive Davenhall To: Bob Mann , Andy Lawrence , "Read, Mike" , Eckhard Sutorius , Mark S. Holliman , Nicholas Cross , Nigel C. Hambly , Rob Blake , Ross Collins , Stelios Voutsinas , Dave Morris , Clive Davenhall Cc: "dante@astro.puc.cl" , James Dunlop , "j.p.emerson@qmul.ac.uk" , ktn keithnoddle.org , "Lorenzo.Rimoldini@unige.ch" , "mcioni@aip.de" , "matt.jarvis@astro.ox.ac.uk" , "mike@ast.cam.ac.uk" , "nigel.metcalfe@durham.ac.uk" , "norman@astro.gla.ac.uk" , "pwl@star.herts.ac.uk" , "rgm@ast.cam.ac.uk" , "Simon.Dye@nottingham.ac.uk" , "s.j.warren@ic.ac.uk" , "tom.shanks@durham.ac.uk" Subject: WFAU VDFS Science Archives progress meeting minutes, 9 September 2016 WFAU VDFS Science Archives progress meeting minutes, 9 September 2016 Present: NJC, ETWS, RSC, RGM, MSH, RPB, ACD Apologies: MAR, STV, DMR, KTN, AL, DoNM: 10:00 am Friday 23rd September 2016 in the Villas Meeting Room MAR in the chair (room is booked). Actions discharged this time: ------------------------------------------------ ACTION: ETWS to transfer SDSS DR12/13. Done. DR13 has been transferred. ACTION: RPB, NJC to fix the errors encountered while trying to make non-survey releases U/14B/UA15, U/14B/UA18 and U/14B/UA19. Discharged partly completed. U/14B/UA15 and U/14B/UA18 have been released, but U/14B/UA19 continues to give problems. However, the action is discharged as, following a further attempt to make non-survey releases, U/14B/UA19 is included in an extended list of releases with problems. See the new action included in 'Non-survey Data Release', below. Actions carried forward from 15/7/2016 meeting: ---------------------------------------------- ACTION: RGM to investigate the size and content (how does it differ from DR9?) of SDSS DR10 to see if it is worth us hosting and whether it is necessary to keep all of the earlier releases. Continues. The first part had been done: the difference between DR9 and DR10 is largely the presence of additional spectra. The part of this action that continues is that RGM had not yet established which early releases could safely be scrapped. RGM also noted that that DR11 and DR12 were now available and now need to be looked at. For further details see the WFAU SDSS strategy wiki page: http://apache.roe.ac.uk/twiki/bin/view/WFAU/SDSSDRPlan We currently have space so it was decided to start transferring DR12/DR13. Perhaps this action is now obsolete? ACTION: ACD to liaise with DMR and STV to establish a set of deadlines for their work. Continues. ACTION: MAR to look at H-ATLAS documentation. Continues. ACTION: NJC to discuss proposed VVVX meeting with RGM and to keep RSC and ACD apprised of developments pertaining to MOONS. Continues. NJC still has to arrange a meeting with RGM. Specific points and new actions: -------------------------------- Project management: * RGM reported continuing problems in keeping UoE Administration's records of staff-to-project allocations up to date with WFAU's actual allocations. However, he emphasised that this was entirely an administrative problem and does not affect the actual projects that staff work on. WFCAM, VISTA and VST updates: * Nothing to report. Networking: * Nothing to report; the networks were operating nominally. WSA/VSA/OSA/GES Operations: * ETWS reported the following progress: - Metadata ingest and JPEG creation for the latest OSA data up to the end of June is complete. - Metadata ingest and JPEG creation for the latest WSA data up to the 1st August is also complete. - VPHAS reseaming is well under way (approx. 55% done). * RPB reported: - The release of VVV DR5 is progressing. Execution of the pawprint code has completed. The ppErrorBits are currently being updated and this seems to be progressing well. - Preparation of the VHS P96 release was in progress and is currently source merging. * NJC added: - Preparation of the Atlas ESO release continues. * Finally RPB noted that at some convenient juncture the various databases would need to be shuffled around the various host machines in order to make optimal use of the hardware bought recently. This rearrangement would be timed to coincide with a lull in the preparation of releases in order to minimise disruption. Hardware and Systems: * Nothing to report. Review of Software Release Deadlines: * Progress towards the software deadlines was reviewed briefly. See: http://apache.roe.ac.uk/twiki/bin/view/WFAU/PrepareSoft Other Software: * ETWS reported that he had: - written a script to migrate data and ingest log files from one server to another, which enables data produced on sahure to be ingested whereas it would otherwise be ignored, - installed a third party software graphics visualisation package on userkaf for RGB's student. It might also be useful for WFAU staff. Review of Survey Release Deadlines: * Progress towards forthcoming survey deadlines was reviewed. See: http://apache.roe.ac.uk/twiki/bin/view/WFAU/ForthcomingDeadlines Survey Data Release: * None. Non-survey Data Release: * RPB reported that the following non-surveys have been successfully released: U/14B/UA15 U/14B/UA18 U/15A/H02 U/15B/H11A1 U/15B/UA02A U/15B/UA08A U/15B/UA11 U/15B/UA12A U/15B/UA13A U/15B/UA17 U/15B/UA18 U/15B/UA19 The following non-surveys have issues: u/14b/ua19 u/15a/na06 u/15a/h11a u/15a/ua14 u/15a/ua09 u/15b/na04 u/15b/h01. ACTION: RPB to check the log files and inform NJC of the problems with the above non-survey releases. ACTION: NJC to investigate and fix the the problems with the above non-survey releases. Astrogrid deployment, VO & Data Analysis services: * MSH is waiting for a usable TAP service from DMR (presumably there's one in the pipeline, but it hasn't come on stream yet. Sorry. Ok, no I'm not). ACTION: MSC to check with DMR about progress towards, and anticipated release date of, a usable TAP service. Miscellaneous: * RPB noted that since the refurbishing of the machine room the the tape library had been unable to talk to userkaf. He had tried muttering various incantations over it, but none had proved efficacious. He now planned to buy a new interface card and to move the tape library to a more recent machine. He added that he was about one third of the way through reading the WFCAM raw data tapes, but could make no further progress until the tape library was working again. He also noted that since the tapes had been returned from storage he'd noticed a small gap in the date sequence of the observations and was unsure whether this was real, corresponding to down-time for the instrument, or meant that some tapes had been misplaced. ACTION: RPB to check with CASU whether they had any WFCAM observations made during the gap in the Edinburgh tapes. ACD, Original: 10/9/16, revised: 15/9/16.